OpenStudy (anonymous):

Help? What is the Least Common Denominator (LCD) of 5 17 -- And -- 16 20 A 5 B 40 C 80 D 85 (Thanks <3)

OpenStudy (anonymous):

80

OpenStudy (anonymous):

16 can't go into any of those except 80. Since 16 and 20 can both go into 80 and it's the smallest number possible, it's 80.

OpenStudy (anonymous):

Thanks guys!!! <3

jimthompson5910 (jim_thompson5910):

Step 1) Find the GCF between 16 and 20 (in this case it's 4) Step 2) Multiply 16 and 20 to get 320 Step 3) Divide that prev result by the GCF 4 to get 320/4 = 80 So the LCM is 80 So the LCD is 80

OpenStudy (anonymous):

\[16=2^4\\ 20 = 2^2 5\\ LCM(16,20)= 2^4 (5) =80 \]

OpenStudy (anonymous):

GCD(16,20)= 2^2=4